### v2

# Preference Optimization with Multi-Sample Comparisons

Abstract
Recent advancements in generative models, particularly large language models (LLMs) and diffusion models, have been driven by extensive pretraining on large datasets followed by post-training. However, current post-training methods such as reinforcement learning from human feedback (RLHF) and direct alignment from preference methods (DAP) primarily utilize single-sample comparisons. These approaches often fail to capture critical characteristics such as generative diversity and bias, which are more accurately assessed through multiple samples. To address these limitations, we introduce a novel approach that extends post-training to include multi-sample comparisons. To achieve this, we propose Multi-sample Direct Preference Optimization (mDPO) and Multi-sample Identity Preference Optimization (mIPO). These methods improve traditional DAP methods by focusing on group-wise characteristics. Empirically, we demonstrate that multi-sample comparison is more effective in optimizing collective characteristics (e.g., diversity and bias) for generative models than single-sample comparison. Additionally, our findings suggest that multi-sample comparisons provide a more robust optimization framework, particularly for dataset with label noise.
[…]
1 Introduction
Generative models, particularly large language models (LLMs) (Achiam et al., 2023; Meta AI, 2024; Bai et al., 2023; Bi et al., 2024; Gemini et al., 2023; Anthropic, 2024) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023), hold the tremendous promise to transform numerous industries by automating complex tasks, enhancing creativity, and personalizing user experiences at an unprecedented scale (Eloundou et al., 2023). These models achieve their capabilities through extensive pretraining on large-scale datasets, followed by post-training to unlock the capabilities (i.e., the superficial alignment hypothesis) (Wei et al., 2022; Tay et al., 2022; Zhou et al., 2024a). For post-training stage, reinforcement learning from human feedback (RLHF) and direct alignment from preference methods (DAP) have been crucial for the success of both LLMs (Rafailov et al., 2024; Azar et al., 2024) and diffusion models (Black et al., 2023; Wallace et al., 2023; Yang et al., 2024).
Figure 1: Top: Diversity of responses from two groups for improving urban transportation. The left group provides a broader range of approaches, including public transit and infrastructure improvements. The right group focuses more narrowly on specific technological and management solutions. Bottom: Bias in images from two groups. The left group displays a more balanced representation of race and gender, while the right group predominantly features males due to stereotypes.
Despite these advancements, current post-training approaches predominantly focus on single-sample comparisons, which fail to capture characteristics better assessed through distributions of samples, such as creativity and bias. Evaluating a model’s creativity/consistency or detecting biases requires analyzing the variability and diversity across multiple outputs, not just individual ones. For example, while LLMs are proficient in crafting narratives, they often show limitations in generating a diverse representation of genres (Patel et al., 2024; Wang et al., 2024). Additionally, these models tend to have lower entropy in their predictive distributions after post-training, leading to limited generative diversity (Mohammadi, 2024; Wang et al., 2023; Wiher et al., 2022; Khalifa et al., 2020).
Consider the task of generating a random integer between 0 and 10. Ideally, the model should not prefer any particular number. However, Zhang et al. (2024a) show that existing models often bias towards certain numbers over the others. Similar issues are observed in diffusion models, which may display biases in the generated outputs based on factors like gender or race (Luccioni et al., 2023; Chen et al., 2024). Inconsistencies in generation is also a crucial issue that needs to be addressed to make models more reliable (Liu et al., 2023; Bubeck et al., 2023). The aforementioned failures cannot be captured by a single sample; instead, they are distributional issues (see Fig. 1 for an illustration). To address the limitations, we extend the post-training paradigm to multi-sample comparisons. This approach involves evaluating the model’s performance over distributions of samples rather than individual samples. By curating groups of responses and assessing their collective characteristics, we can better align the model’s outputs with desired distributional properties.
In this work, we introduce Multi-sample Direct Preference Optimization (mDPO) and Multi-sample Identity Preference Optimization (mIPO), which are extensions of the prior DAP methods–DPO (Rafailov et al., 2024) and IPO (Azar et al., 2024) [^1] Our framework can also be extended to other preference optimization algorithms.. Unlike their predecessors, which rely on single-sample comparisons, mDPO and mIPO utilize multi-sample comparisons to better capture group-wise or distributional characteristics. Our experiments involve fine-tuning language models to generate random numbers in a calibrated manner and enhancing the diversity of genres in creative story writing. Additionally, for diffusion models, we demonstrate a significant reduction in gender and race biases compared to the traditional single-sample comparison approach. Furthermore, we present remarkable results where mDPO and mIPO can iteratively improve the model quality by only contrasting multiple responses generated by a stronger and weaker model without using precise pairwise comparisons.
2 Related Works
Reinforcement Learning from Human Feedback (RLHF) has been pivotal in advancing the capabilities of large language models (Ouyang et al., 2022; Bai et al., 2022) and, more broadly, in agent alignment (Christiano et al., 2017; Leike et al., 2018). RLHF was initially introduced by Christiano et al. (2017) to tackle classic reinforcement learning tasks such as those found in MuJoCo. The work by Ouyang et al. (2022) marked the first significant application of RLHF in aligning language models to follow human instructions, thereby establishing a foundation for subsequent models like ChatGPT (Achiam et al., 2023). To address the complexities and instabilities associated with RL methods, Direct Alignment from Preference (DAP) methods (Rafailov et al., 2024; Azar et al., 2024; Dong et al., 2023; Yuan et al., 2023; Zhao et al., 2023) have been proposed. Notably, Rafailov et al. (2024) introduced Direct Preference Optimization (DPO), which reframes the RL problem into a supervised learning problem by deriving an analytical relationship between policy and reward. To counteract potential overoptimization issues in DPO, Azar et al. (2024) proposed Identity Preference Optimization (IPO), which employs an $\ell_{2}$ loss to regress the margin towards a predefined threshold. In the context of diffusion models, Black et al. (2023) introduced Denoising Diffusion Policy Optimization (DDPO) to optimize diffusion models using specific reward functions. Wallace et al. (2023) expanded the application of DPO from language domains to image domains. However, these methods predominantly focus on single-sample comparisons and may not adequately capture the collective characteristics of output distributions.
Distributional difference (Zhong et al., 2022) describes the difference between two or more distributions, which is a generalization of single-sample comparison. While many characteristics can be judged from a single sample, some properties can only be deduced from multiple samples, such as diversity and bias (Santurkar et al., 2023; Chen et al., 2024; Zhang et al., 2024b; Zhou et al., 2024b; Mohammadi, 2024; Wang et al., 2023; Go et al., 2023). To measure the distributional difference, Zhong et al. (2022) proposed a hypothesis-verifier framework to summarize the difference between two corpora. Dunlap et al. (2024) further extended this framework to the image domain to capture the difference between two sets of images. More recently, Zhong et al. (2023) considered the problem of distributional difference in a goal-driven setting. Melnyk et al. (2024) addressed the problem of distributional alignment for language models using optimal transport, aiming to optimize the chosen responses across all prompts jointly. In contrast, our work addresses an orthogonal problem by focusing on optimizing the distributional preference under the same prompt.
While finalizing this manuscript for submission to arXiv, we became aware of a related study by Li et al. (2024). Both works independently address the formulation of multi-sample comparison. Our contributions differ in several key aspects, such as the introduction of additional algorithms, distinct experimental foci, and differing applications. Notably, our primary focus is on language models, with an emphasis on providing a low-variance and potentially unbiased estimator. In contrast, their work centers around text-to-image tasks and does not include the IPO-based methods we derived.
3 Preliminaries
Supervised Finetuning. After the language model has been pretrained extensively on large corpora, the next step is typically supervised finetuning (SFT). This process aims to teach the model the dialog formats and refine its predictions to interact more naturally with human and prepare it for the subsequent alignment stages. Specifically, the pretrained language model is finetuned using a supervised dataset $\mathcal{D}=\{(x,y)_{i}\}_{i=1}^{N}$, where $x$ represents the input and $y$ denotes the target. The objective of SFT is to minimize the negative log-likelihood,
$\mathcal{L}_{\text{SFT}}(\pi_{{\bm{\theta}}},\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[-\log\pi_{{\bm{\theta}}}(y|x)\right].$
Direct Alignment from Preference.
For single-sample comparison setting, the dataset consists of $\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$, where $x$ denotes the prompt or the input to the model, $y_{w}$ is the preferred response over $y_{l}$. Under the Bradley-Terry model, the likelihood of $y_{w}$ is preferred to $y_{l}$ is computed as
$p(y_{w}\succcurlyeq y_{l}|x)=\sigma(r(y_{w},x)-r(y_{l},x))=\frac{\exp(r(y_{w},x))}{\exp(r(y_{w},x))+\exp(r(y_{l},x))},$
where $r(y,x)$ is the scalar reward of answering $y$ given $x$. Direct preference optimization (DPO) (Rafailov et al., 2024) establishes an analytical form between the reward function and the policy language model, where the language model’s prediction represents an implicit reward. Specifically, under DPO, the implicit reward can be computed by
$r_{\bm{\theta}}(y,x)=\beta\log\frac{\pi_{\bm{\theta}}(y|x)}{\pi_{\textnormal{ref}}(y|x)}+\mathrm{const}(x).$
Using this implicit reward, DPO optimizes the policy $\pi_{\bm{\theta}}$ by minimizing the negative log-likelihood on the offline preference dataset,
$\mathcal{L}_{\mathrm{DPO}}(\pi_{\bm{\theta}},\mathcal{D})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\bm{\theta}}(y_{w},x)-r_{\bm{\theta}}(y_{l},x)\right)\right].$
The Bradley-Terry model is based on the assumption of Gumbel noise in the reward function, which is well-suited for discrete random variables. However, it has certain limitations such as overfitting, as noted by Azar et al. (2024). Specifically, Identity Preference Optimization (IPO) (Azar et al., 2024) addresses these limitations by bypassing the Bradley-Terry model for preference modeling. IPO aims to mitigate the issue of “overfitting" in preference datasets by replacing the sigmoid function with the squared distance,
$\mathcal{L}_{\mathrm{IPO}}(\pi_{\bm{\theta}},\mathcal{D})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg{[}\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\frac{\tau^{-1}}{2}\bigg{]}^{2},$
which can be interpreted as regressing the difference between the log-likelihood ratio to $\tau^{-1}/2$.
4 Method
The singleton preference (i.e., the marginal distribution) may not generalize well to the distributional preference (i.e., the joint distribution) (Melnyk et al., 2024). For most of the cases, when given a prompt $x$, there is indeed a preference between two responses $y_{1}$ and $y_{2}$. However, there are also cases where the preference cannot be assigned purely based on two singletons. Consider the cases where we prompt the diffusion models to generate an image of “a software engineer”. For such cases, there is no preference between “a female software engineer” and “a male software engineer”. Similarly, if a language model is prompted to generate an integer randomly and uniformly from $\{0,1,...,9\}$, there is also no preference between generating any two specific numbers, e.g., 3 or 5. Nonetheless, preferences can exist at a multi-sample or joint distribution level.
4.1 Multi-sample Preference Optimization
We generalize the preference model from singleton comparison to multi-sample comparison, where the preference is assigned at the multi-sample level instead of on a singleton-level. Let $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$ denote two group of samples drawn from some models’ conditional distributions when the prompt $x$ is given. [^2] The definition of a group $\mathcal{G}$ can be generalized to distributions. Thus, $\mathcal{G}$ is equivalent to the predictive distribution of the language model. Following the definitions from Azar et al. (2024), we define the likelihood of $\mathcal{G}_{w}$ preferred over $\mathcal{G}_{l}$ as
$p(\mathcal{G}_{w}\succcurlyeq\mathcal{G}_{l}|x)=\Phi(r(\mathcal{G}_{w},x)-r(\mathcal{G}_{l},x)),$
where $\Phi$ can be a sigmoid function (recovering the Bradley-Terry model). Given the reward function $r(\cdot,\cdot)$, the goal of RLHF is to maximize the reward under reverse KL constraints. This can be captured by the following optimization objective,
$\max_{{\bm{\theta}}}\mathbb{E}_{\mathcal{G}\sim\pi_{\bm{\theta}}|x,x\sim p_{x}}\left[r(\mathcal{G},x)-\beta\cdot\text{KL}(\pi_{\bm{\theta}}(\cdot|x)||\pi_{\text{ref}}(\cdot|x))\right].$
Generalizing the derivation of DPO, we obtain the following equation for the reward function,
$r(\mathcal{G},x)=\beta\log\frac{\pi_{\bm{\theta}}(\mathcal{G}|x)}{\pi_{\textnormal{ref}}(\mathcal{G}|x)}+\mathrm{const}(x).$
where $\text{const}(x)$ is some constant that only depends $x$, and $\beta$ is the coefficient of the KL regularization. For a given dataset, $\{(\mathcal{G}_{w},\mathcal{G}_{l},x)_{i}\}_{i=1}^{N}$, where $\mathcal{G}=\{y_{j}\}_{j=1}^{k}$ with $y_{j}\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\pi(\cdot|x)$, we can expand the likelihood by $\pi_{\theta}(\mathcal{G}|x)=\prod_{y\in\mathcal{G}}\pi_{\bm{\theta}}(y|x)$. To avoid the effect of the size of $\mathcal{G}$, we consider using the geometric mean instead, $\pi_{\theta}(\mathcal{G}|x)=(\prod_{y\in\mathcal{G}}\pi_{\bm{\theta}}(y|x))^{1/|\mathcal{G}|}$. Thus the objective of multi-sample DPO becomes (where $\Phi(z)=\sigma(z)$),
$\mathcal{L}_{\mathrm{mDPO}}=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\bigg{[}-\log\sigma\bigg{(}\beta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-\beta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]\bigg{)}\bigg{]}.$ (1)
Similarly, for the IPO variant (where $\Phi(z)=(z-\tau^{-1}/2)^{2}$), we have
$\mathcal{L}_{\mathrm{mIPO}}=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\bigg{[}\bigg{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]-\frac{\tau^{-1}}{2}\bigg{)}^{2}\bigg{]}.$ (2)
The main difference (highlighted in Eq. (1) and Eq. (2) with red) from their original objectives is that the implicit reward for a single sample is replaced with the expectation of the implicit reward from the distribution $\mathcal{G}_{w}$ or $\mathcal{G}_{l}$. In practice, we use a finite set of examples collected from $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$ and take the average of their implicit rewards to approximate the expectations. In the next section, we analyze the bias and variance of this approximation.
4.2 Stochastic Estimator for Efficient Optimization
Figure 2: Biased estimator vs. Unbiased estimator.
Our focus is on finetuning large generative models, and thus a scalable estimator for the gradient / objective is necessary to make the algorithm practically useful. Deriving a mini-batch (potentially unbiased and low-variance) estimator for mDPO is challenging due to the non-linearity of the sigmoid function. However, for mIPO, this task is more tractable by expanding the square function. This is equivalent to computing the unbiased estimator for the following objective, $\ell=\left(\mathbb{E}_{x\sim p}[f(x)]-\mathbb{E}_{x\sim q}[f(x)]-c\right)^{2}.$ The unbiased estimator for mIPO is stated in the following result. {restatable}propunbiasedestimator Let $f:\mathcal{X}\to\mathbb{R}$ be a measurable function, and let $p$ and $q$ be probability distributions on $\mathcal{X}$. Define $\ell=(\mathbb{E}_{x\sim p}[f(x)]-\mathbb{E}_{x\sim q}[f(x)]-c)^{2}$, where $c$ is a constant. Let $x^{p}_{1},\ldots,x^{p}_{n}$ be i.i.d. samples from $p$, and $x^{q}_{1},\ldots,x^{q}_{m}$ be i.i.d. samples from $q$. Then,
^$\hat{\ell}=\left(\frac{1}{n}\sum_{i=1}^{n}f(x^{p}_{i})-\frac{1}{m}\sum_{j=1}^{m}f(x^{q}_{j})-c\right)^{2}-\left(\frac{\hat{\sigma}_{p}^{2}}{n}+\frac{\hat{\sigma}_{q}^{2}}{m}\right)$
is an unbiased estimator of $\ell$, where $\hat{\sigma}_{p}^{2}$ and $\hat{\sigma}_{q}^{2}$ are the sample variances of $f(x)$ under $p$ and $q$. Therefore, to optimize the mIPO objective, we can sample mini-batches similar to IPO. Each mini-batch can consist of multiple responses from both $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$. While increasing the number of samples from each group reduces the variance of the estimated gradient, it also raises computational and memory costs. In summary, the mIPO objective is thus
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\Bigg{[}\Bigg{(}\frac{1}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{1}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}-\frac{\tau^{-1}}{2}\Bigg{)}^{2}-\frac{\hat{\sigma}_{w}^{2}}{k}-\frac{\hat{\sigma}_{l}^{2}}{k}\Bigg{]}.$
For mDPO, we consider the following objective for multi-sample comparison, which may exhibit low variance (for larger values of $k$) but is biased,
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\Bigg{[}-\log\sigma\Bigg{(}\frac{\beta}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{\beta}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}\Bigg{)}\Bigg{]}.$
To further understand the variance, we present the following Proposition 4.2, which shows that the variance of the estimator $\hat{\ell}$ is $\tilde{\mathcal{O}}(1/k)$, where $k$ is the number of samples. {restatable}propvarianceofestimator Let $\mu_{p}=\mathbb{E}_{x\sim p}[f(x)]$, $\mu_{q}=\mathbb{E}_{x\sim q}[f(x)]$, $\sigma_{p}^{2}=\text{Var}_{x\sim p}[f(x)]$, $\sigma_{q}^{2}=\text{Var}_{x\sim q}[f(x)]$ and $n$ and $m$ be the number of independent samples from distributions $p$ and $q$, respectively. Then, the variance of the mini-batch estimator $\hat{\ell}$ is given by
$\text{Var}(\hat{\ell})=\mathcal{O}\left(\left(\frac{\sigma_{p}^{2}}{n}+\frac{\sigma_{q}^{2}}{m}\right)\cdot\left(\frac{\sigma_{p}^{2}}{n}+\frac{\sigma_{q}^{2}}{m}+(\mu_{p}-\mu_{q}-c)^{2}\right)\right).$
To gain an understanding about the variance and bias, we run a simulation using the function $f(x)=x^{2}$ to assess the effects of variance and bias and their relationship to the sample size $k$. The trend is plotted in Fig. 2. We observe that the unbiased estimator results in lower error compared to the biased estimator at small sample sizes. However, as the sample size increases, the error of the biased estimator also decreases and performs similarly to the unbiased one. This implies that as the sample size increases, the choice of estimator may become less critical.
5 Experiments
5.1 Random Number Generator (RNG)
Figure 3: Distance comparison.
Despite existing LLMs being trained to follow instructions and perform well on many tasks, they perform surprisingly poorly when asked to generate random outputs (See Fig. 3). A canonical example is instructing LLMs to generate a random number within a given interval. Existing LLMs exhibit a strong bias towards certain numbers rather than producing purely random outputs. In such settings, single-sample comparisons may be ineffective for modeling the distributional properties of the model’s outputs, necessitating the use of multi-sample comparisons.
Preference data construction and finetuning. Our initial experiments focus on generating integers uniformly at random within an interval $[a,b]$, where $a$ and $b$ are integers. To create the dataset, we sample $a$ randomly from $[0,1000]$ and the interval gap $m$ from $[5,10]$, with $b$ defined as $a+m$. The preferred response group is based on a uniform distribution of numbers, while non-uniform distributions are classified as rejected. We generated approximately 3,000 paired groups for training. For testing, we sample $a$ and $b$ from $[0,1000]$ with $a<b$ but do not impose the same gap constraints as in the training set to evaluate generalizability. We implemented the multi-sample versions of IPO and DPO, named mIPO and mDPO, respectively. To stabilize training, we add a negative log-likelihood [^3] Without this, the language model tends to generate numbers beyond the specified range. We discuss this further from a constrained optimization perspective. to the objective of both the multi-sample and original versions of IPO and DPO, shown to be effective in reasoning tasks (Pang et al., 2024). For finetuning, we use the Llama 3-8B instruct model with LoRA (r=64, $\alpha=128$). More details are in Appendix 7.1.
Metrics and results. We compare the predictive entropy for each model with 100 testing prompts instructing the model to generate random numbers within specific intervals. We calculate the average win rates for the same prompt, with the winner being the one with larger entropy. The results are presented in Table 1. We observe that mIPO, IPO, mDPO, and DPO consistently and significantly outperform the SFT baseline. Additionally, both mIPO and mDPO outperform IPO and DPO by a large margin. Figure 3 shows the predictive distribution before and after finetuning. The original Llama3 model significantly favors the number 7 over others. However, after finetuning, the predictive distribution is much closer to the uniform distribution.
Table 1: Comparison of Win Rates in the Random Number Generation Experiment with Llama-3-8B: Win rates are calculated based on the uniformity of the distribution given the prompt.
mIPO vs IPO mIPO vs SFT IPO vs SFT mDPO vs DPO mDPO vs SFT DPO vs SFT
Win Rate 0.95 0.99 0.98 0.80 0.99 0.99
5.2 Controlled Debiasing for Image Generation
Diffusion models (Sohl-Dickstein et al., 2015) have emerged as a powerful workhorse for image generation (Ho et al., 2020; Rombach et al., 2022; Dai et al., 2023; Esser et al., 2024). Despite their success, the generated images often reflect strong biases and stereotypes due to imbalances in the training data (Ananya, 2024). These biases can result in harmful societal stereotypes and limit the diversity of generated content. To illustrate this point, we visualize the distribution of generated images for different occupations and highlight the discrimination in race and biases in Figure 10 in the Appendix using Stable diffusion 1.5 (Rombach et al., 2022). To address the generation biases, we extend the diffusion DPO objective proposed in Wallace et al. (2023) from single-sample comparison to multi-sample comparison in a similar way as we discussed in Section 4. For diffusion models, the mDPO objective can be formulated as [^4] It’s more appropriate to interpret $r(\mathbf{\epsilon},{\bm{x}};\theta)$ as a cost rather than a reward. Consequently, the formulation uses $-\beta$ instead of $\beta$, as done in the original DPO.
$\mathcal{L}_{\text{mDPO}}^{\text{diff}}(\epsilon_{\bm{\theta}},\mathcal{D})=\mathbb{E}\big{[}-\log\sigma\big{(}-\beta T\omega(\lambda_{t})\cdot\big{(}\mathbb{E}_{\mathcal{G}_{w}}\left[r(\mathbf{\epsilon}^{w},{\bm{x}}_{t}^{w},t;{\bm{\theta}})\right]-\mathbb{E}_{\mathcal{G}_{l}}\left[r(\mathbf{\epsilon}^{l},{\bm{x}}_{t}^{l},t;{\bm{\theta}})\right]\big{)}\big{)}\big{]},$ (3)
where $r(\mathbf{\epsilon}^{w},{\bm{x}}_{t}^{w},t;{\bm{\theta}})=\|\mathbf{\epsilon}^{w}-\mathbf{\epsilon}_{\bm{\theta}}({\bm{x}}^{w}_{t},t;c)\|_{2}^{2}-\|\mathbf{\epsilon}^{w}-\mathbf{\epsilon}_{\mathrm{ref}}({\bm{x}}^{w}_{t},t;c)\|_{2}^{2}$, $\epsilon^{w}={\bm{x}}^{w}_{t}-{\bm{x}}^{w}$ and ${\bm{x}}^{w}\sim\mathcal{G}_{w}$, and the expectation is taken over $(\mathcal{G}_{w},\mathcal{G}_{l},c)\sim\mathcal{D}$ and $t\sim\mathcal{U}(0,T)$. The same applies to $\mathcal{G}_{l}$, ${\bm{x}}_{t}^{l}$ and ${\bm{x}}^{l}$. We adopt the Stable Diffusion 1.5 (Rombach et al., 2022) as our base model due to its popularity within the community, which we also find has strong bias in favor of certain genders and races.
Dataset construction and finetuning. To construct the dataset, we first collect $50$ occupations, and take $80\%$ of the occupations for generating training data and the remaining $20\%$ for evaluation. For each occupation, we generate $6$ images with varied races in {black, white, aisan} and genders in {male, female}, which form our chosen group. For the rejected group, we use the images generated from the original Stable diffusion 1.5 via API [^5] https://stablediffusionapi.com/docs/category/stable-diffusion-api. Our training code is adapted from the original implementation by Wallace et al. (2023). More details on the training can be found in Appendix 7.2.
Evaluation metric and results. To measure the generation quality, we used the Simpson Diversity Index (Simpson, 1949), which focuses more on the dominance and even distribution of species. To compute it, we use $D=1-\sum_{i}(n_{i}/N)^{2}$, where $n_{i}$ is the number of instance falls in $i_{\text{th}}$ category (e.g., for gender, the categories we consider are {male, female}), and $N=\sum_{i}n_{i}$.
Table 2: Averaged Simpson Diversity Index for the generated images of different occupations.
Original SFT DPO mDPO
Gender $\uparrow$ 0.110 0.318 0.283 0.353
Race $\uparrow$ 0.124 0.454 0.447 0.516
In general, the higher the value, the better the diversity. We compared three methods, SFT, DPO (equivalent to $k=1$ for mDPO), and mDPO ($k=6$). The results are presented in Table 2. We observe that mDPO performs the best in both metrics, improving the Simpson Diversity Index significantly over the original SD 1.5. For qualitative comparison, we visualized $10$ images generated by the original diffusion model and the diffusion model after mDPO finetuning in Appendix 8.1 using the testing prompt. We observe that the finetuned model generates more balanced and diversified images in terms of race and gender than the baseline.
5.3 Improving Quality of Creative Fiction Generation
Fiction generation represents a crucial aspect of the broader field of creative writing, and serves as an essential benchmark for evaluating the capabilities of LLMs (Gómez-Rodríguez and Williams, 2023; Mohammadi, 2024). Despite the proficiency of current LLMs in crafting creative narratives, a significant challenge remains in ensuring a diverse representation of genres (Patel et al., 2024; Wang et al., 2024) while maintaining high writing quality. Ideally, when prompted with a consistent fiction topic, LLMs should be capable of generating distinctly different stories across various genres, closely aligning with user intentions, even in the absence of explicit genre specifications in the prompt. This diversity should extend beyond mere lexical variations, incorporating unique genre-specific elements in each narrative output. In this section, we assess whether the proposed mDPO and mIPO can help fine-tune LLMs to achieve higher quality and more diverse fiction generations. Similar to the previous sections, we compare mDPO and mIPO, with their original baselines.
Preference data construction and finetuning. We utilized over 8,000 publically available prompts [^6] https://draftsparks.com/browse/fiction-prompts/ for fiction generation in constructing our preference dataset. Each prompt was submitted five times to both the Llama 2-7B and Llama 3-8B models. For prompts to Llama 3-8B, we explicitly defined the genres including fantasy, sci-fi, mystery, romance, and horror, to enhance genre diversity, while keeping the prompts for Llama 2-7B unchanged. The responses from Llama 3-8B comprised the chosen set in our preference dataset, whereas the responses from Llama 2-7B (Touvron et al., 2023) formed the rejected set. When finetuning with the single-sample DPO and IPO baselines, one response from each chosen and rejected set was selected. Conversely, with mDPO and mIPO, $k$ responses were selected accordingly. We finetuned the Llama 3-8B model using this preference dataset. More details, including finetuning parameters and specific prompts used during data generation, are illustrated in Appendix 7.3.1 and 7.3.2.
Evaluation metrics and results.
Figure 4: Diversity in fiction generation using the same model (Llama 3-8B) finetuned with different approaches, assessed through genre distribution. Left: mDPO and DPO. Right: mIPO and IPO. The KL-divergences between different genre distributions and the uniform distribution are (smaller is better, and the best ones are highlighted in bold font.) DPO: $0.170$; mDPO ($k=3$): $\mathbf{0.126}$; mDPO ($k=5$): $0.142$; IPO: $0.125$; mIPO ($k=3$): $0.094$; mIPO ($k=5$): $\mathbf{0.050}$.
Figure 5: mDPO and mIPO versus DPO and IPO on Alpaca Evals using GPT-4o evaluation.
We primarily evaluated the fine-tuned Llama 3-8B using two metrics: writing quality and genre diversity. For assessing writing quality, we employed the evaluation rubrics proposed by Chakrabarty et al. (2024), which measure fiction quality across four dimensions: fluency, flexibility, originality, and elaboration. Each dimension is specifically evaluated using a total of 14 Yes/No questions. To measure the diversity, we measure entropy of the generated genre diversity and lexical diversity, such as distinct-n (Li et al., 2015). In our experiment, we combined all the questions and used GPT-4o (Achiam et al., 2023) as the judge. Each “Yes" was scored as 1 and each “No" as 0, with the maximum possible score for a generated fiction being 14. We then compared the final scores of Llama 3-8B finetuned with mDPO, mIPO, and their single-sample baselines.
Specific evaluation prompts are shown in Appendix 7.3.3. We assess genre diversity using both lexical-level metrics, including the number of distinct unigrams (distinct-1) and bigrams (distinct-2) (Li et al., 2015), as well as semantic-level genre distribution identified by GPT-4o. Details regarding the prompt used for genre identification are provided in Appendix 7.3.4.
Table 3: Quality comparison of creative fiction writing. The maximum score a model can achieve is 14.
DPO mDPO $(k=3)$ mDPO $(k=5)$ IPO mIPO $(k=3)$ mIPO $(k=5)$
Quality 10.570 10.671 11.483 10.623 11.190 10.806
The results of the fiction writing quality evaluations are presented in Table 3. It is evident that the proposed multi-sample approaches have a clear advantage over the single-sample baselines, despite the more comprehensive and well-rounded challenge of enhancing the creative writing capabilities (Mohammadi, 2024). The diversity evaluation results, based on the number of unique unigrams and bigrams as well as genre distributions, are shown in Table 4 and Fig. 4, respectively.
Table 4: Comparison of lexical-level diversity between the proposed mDPO, mIPO, and baseline methods.
DPO mDPO $(k=3)$ mDPO $(k=5)$ IPO mIPO $(k=3)$ mIPO $(k=5)$
distinct-1 $\uparrow$ 0.025 0.026 0.027 0.025 0.026 0.032
distinct-2 $\uparrow$ 0.187 0.189 0.192 0.189 0.185 0.209
To better illustrate the differences in diversity in fiction generation, we calculate the KL-divergence between the distributions shown in Fig. 4 and the uniform distribution, with the results presented in the caption. We observe that models finetuned using multi-sample methods exhibit improved diversity at both lexical and semantic levels compared to the baselines.
5.4 Training with Llama3-70B vs. Llama3-8B Generated Preference Data
Synthetic data is becoming increasingly prevalent due to its scalability and lower cost compared to human labeling (Meta AI, 2024; Adler et al., 2024; Liu et al., 2024). Our last set of experiments aims to demonstrate the efficacy of our method in handling synthetic datasets with inherent label noise. This is particularly useful in iterative alignment scenarios, where we train the model iteratively to enhance alignment performance. In these cases, we might have two qualitatively good models from prior iterations, but one may outperform the other in average. However, when examining individual responses, the performance of these models may not always be consistently better than the other. In such situations, mDPO or mIPO algorithms should be more robust than DPO or IPO. This intuition is supported by the following remarks and subsequent experiments.
{restatable}
remmeanvssingle For two independent and bounded random variables $X$ and $Y$, if $\mathbb{E}[X]-\mathbb{E}[Y]>0$, then the probability $p(\sum_{i=1}^{k}X_{i}>\sum_{i=1}^{k}Y_{i})$ will (approximately) increase as the sample size $k$ increases. Therefore, the multi-sample pairwise comparison is (approximately) more likely to be correct than the single-sample pairwise comparison. In the asymptotic setting ($k\uparrow\infty$), the probability will converge to $1$ as $\mathbb{E}[X]>\mathbb{E}[Y]$.
Figure 6: Left: The impact of $k$ for mDPO evaluated using GPT-4o; Middle: The impact of $k$ for mIPO evaluated using GPT-4o; and Right: Iterative improvement with mDPO.
Figure 7: Evaluation of multi- and single-sample comparison under varying label conditions. The left two plots (green) depict performance in a noise-free setting using GPT-4o labels, the middle two plots (red) show results with default (noisy) labeling, and the right three plots (blue) compare the performance gap between noise-free and noisy settings.
To empirically validate our intuition and demonstrate the effectiveness of our method, we conduct experiments using the Alpaca benchmark (Dubois et al., 2024) with the Llama 3-8B base model (Meta AI, 2024). Initially, we use the instruct versions of Llama 3-8B and Llama 3-70B to generate five responses for each prompt [^7] We used outputs generated by Llama models instead of the original outputs from the Alpaca dataset.. We then select the responses generated by the 70B model as the chosen group and those from the 7B model as the rejected group. We follow the Alpaca training procedures for SFT and RLHF. We first finetune the base model on the Alpaca SFT dataset, and then apply mDPO or mIPO with $k=1,2,...,5$ on the synthetic data.
Results with varying $k$. The results for varying group sizes $k$ are shown in Fig. 5, where win rates are computed against models trained using DPO and IPO with the same dataset. The results indicate that both mDPO and mIPO significantly outperform the baselines. Additionally, we perform an ablation study to examine the effect of $k$ using mDPO and mIPO. As illustrated in the leftmost and middle plots of Fig. 6, we observe that as the value of $k$ increases, the win rates improve while the lose rates decrease for most values of $k$, except for $k=5$ for mIPO.
Results on iterative improvement. Lastly, we conduct experiments to demonstrate the effectiveness of our method in iteratively improving the alignment performance. We use data generated from previous rounds to form the preference data and then apply our method with $k=5$ for iterative fine-tuning of the model. The results, shown in the rightmost plot of Fig. 6, reveal that with each iteration, our method consistently enhances the win rate compared to the baseline in terms win rates.
Choice between mDPO and DPO? We used GPT-4o to label pairs of sample groups generated by Llama 3-8B and 70B for $k=1,2,3$ to simulate a noise-free setting. To reduce GPT-4o labeling costs, we sampled 20% of the original synthetic dataset. We refer to the dataset with the original preference label as the default label and the GPT-4o-generated label as the GPT-4o label. Our experiments reveal that without labeling noise, multi-sample comparison has no advantage over single-sample comparison, as shown in the left two green plots in Fig. 7. These plots display the win rates for $k=2$ and $k=3$ compared to $k=1$, with $k=2$ showing a slight improvement over $k=1$. To confirm the impact of potentially noisy labels, we switched to the default labeling. In this case, both $k=2$ and $k=3$ outperformed $k=1$, as shown in the middle two red plots in Fig. 7, consistent with our prior results. Lastly, to quantify effect of label noise, we computed the win rate of models trained using default labeling versus GPT-4o labeling for $k=1,2,3$. The right three blue plots in Fig. 7 demonstrate that default labeling performs significantly worse than GPT-4o labeling. However, as $k$ increases, the gap between win and lose rates narrows, confirming that multi-sample comparison is more robust to labeling noise. In conclusion, multi-sample comparison is much more advantageous with labeling noise, while single-sample comparison is best with noise-free labels.
6 Conclusions
In this paper, we introduced multi-sample DPO (mDPO) and multi-sample IPO (mIPO), which are novel extensions to the existing DAP methods. By leveraging multi-sample comparisons, mDPO and mIPO address the limitations of traditional single-sample approaches, offering a more robust framework for optimizing collective characteristics such as diversity and bias in generative models. Comprehensive empirical studies demonstrate the effectiveness of the proposed methods across various domains. In random number generation (§5.1), both mDPO and mIPO achieve higher uniformity. In text-to-image generation (§5.2), multi-sample optimization enables notable reductions in gender and race biases, enhancing the overall fairness and representation in generated images. In creative fiction generation (§5.3), both the quality and diversity of outputs are significantly improved. We further demonstrate that the proposed methods are especially robust against label noise (§5.4).